OpenStudy (anonymous):

Find the integral of (tan(theta)-1)^2 d*theta.

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

\[\int (\tan\theta-1)^2~d\theta~~?\] Expand the binomial: \[(\tan\theta-1)^2=\color{red}{\tan^2\theta}-2\tan\theta\color{red}{+1}=\color{red}{\sec^2\theta}-2\tan\theta\] (red indicates the use of an identity. \[\int\sec^2\theta~d\theta-2\int\tan\theta~d\theta\\ \int\sec^2\theta~d\theta-2\int\frac{\sin\theta}{\cos\theta}~d\theta\] First is easy, the second requires a substitution.